OpenStudy (anonymous):

A company estimates that it will sell N(x) hair dryers after spending $x thousand on advertising as given by: N(x)=-0.25x^4+13x^3-180x^2+10,000 for 15<=x<=24 When is the rate of change of sales increasing, decreasing?(Hint: Find N'(x) which is the rate of change, then find N"(x) to check when N'(x) increasing or decreasing)

jimthompson5910 (jim_thompson5910):

If N(x) = -0.25x^4+13x^3-180x^2+10000 then N ' (x) = ???

OpenStudy (anonymous):

-1x^3+39x^2-180

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now what is N '' (x)

OpenStudy (anonymous):

-3x^2+78x

jimthompson5910 (jim_thompson5910):

very nice

OpenStudy (anonymous):

Im almost a pro haha

jimthompson5910 (jim_thompson5910):

you're definitely a pro

jimthompson5910 (jim_thompson5910):

Now we need to find the roots of N '' (x) = -3x^2+78x N '' (x) = -3x^2+78x 0 = -3x^2+78x -3x^2+78x = 0 -3(x-26) = 0 x - 26 = 0 x = 26 So x = 26 is a possible point of inflection

jimthompson5910 (jim_thompson5910):

If x = 25 (a value to the left of x = 26), then N '' (x) = -3x^2+78x N '' (25) = -3(25)^2+78(25) N '' (25) = 75 This tells us that the rate of change is increasing on the interval (-infinity, 26)

jimthompson5910 (jim_thompson5910):

If x = 27, then N '' (x) = -3x^2+78x N '' (27) = -3(27)^2+78(27) N '' (27) = -81 which means that the rate of change is now decreasing on the interval (26, infinity)

jimthompson5910 (jim_thompson5910):

if we restrict the interval to 15 <= x <= 24 (just saw this part), then the rate of change of the sales is increasing because 15 <= x <= 24 is definitely in the interval (-infinity, 26)

OpenStudy (anonymous):

okay, so the increasing interval is set

jimthompson5910 (jim_thompson5910):

since the derivative is increasing throughout the interval 15 <= x <= 24, there are no intervals where it is decreasing

jimthompson5910 (jim_thompson5910):

so I would say NA or none or something like that

OpenStudy (anonymous):

oh wow that was a lot easier than I thought!

jimthompson5910 (jim_thompson5910):

yeah not too bad

jimthompson5910 (jim_thompson5910):

this is only true because 15 <= x <= 24 is all below 26 where everything is increasing if you stretched it from 15 <= x <= 24 to 15 <= x <= 34, then it would be a different story

OpenStudy (anonymous):

I want to know why n'(x) isn't -1x^3+39x^2-360x?